# Optimal Transport for $\epsilon$-Contaminated Credal Sets

### v1

@@ 3 Main Results
In this work, we give interesting results pertaining a special instance of the core of a lower probability. In particular, we consider $\epsilon$-contaminated credal sets. That is, given a countably additive probability measure $P$ on $\mathcal{X}$, we consider the set
$\mathcal{P}_{\epsilon}=\left\{{\Pi\in\Delta_{\mathcal{X}}:\Pi(A)=(1-\epsilon)P(A)+\epsilon R(A),\quad\forall R\in\Delta_{\mathcal{X}},\forall A\in\mathcal{F}}\right\},$ (4)
where $\epsilon$ is a parameter in $(0,1)$.
@@
Let $\mathcal{P}_{\epsilon}$ be an $\epsilon$-contaminated credal set as in (4). Then,
$\underline{P}^{\prime}(A)=\inf_{\Pi\in\mathcal{P}_{\epsilon}}\Pi(A)=\begin{cases}(1-\epsilon)P(A),&\text{for all }A\in\mathcal{F}\setminus\mathcal{X}\\ 1,&\text{for }A=\mathcal{X}\end{cases}$ (5)
@@
$\mathcal{P}_{\epsilon}=\mathcal{M}(\underline{P}^{\prime})=\{\Pi\in\Delta_{\mathcal{X}}:\Pi(A)\geq\underline{P}^{\prime}(A)\text{, }\forall A\in\mathcal{F}\}.$
@@
Both these properties were proven in Wasserman and Kadane (1990, Example 3). ∎
In the remainder of the paper, we will work with lower probability $\underline{P}$ such that
@@
in place of $\underline{P}^{\prime}$. The reason is twofold: calculations are easier to carry out with $\underline{P}$, and also the following holds.
@@
Pick any countably additive probability measure $P$ on $\mathcal{X}$, any $\epsilon\in(0,1)$, and consider the two lower probabilities $\underline{P}^{\prime}$ and $\underline{P}$ in (5) and (6), respectively. Let $\mathcal{M}(\underline{P})=\{\Pi\in\Delta_{\mathcal{X}}:\Pi(A)\geq\underline{P}(A)\text{, }\forall A\in\mathcal{F}\}$. Then, $\mathcal{M}(\underline{P}^{\prime})=\mathcal{M}(\underline{P})$.
@@
We begin by noting that $\underline{P}^{\prime}$ is a well-defined lower probability by Cerreia-Vioglio et al. (2015, Section 2.1.(viii)).
Now, pick any $\check{\Pi}\in\mathcal{M}(\underline{P}^{\prime})$. By definition, we have that $\check{\Pi}(A)\geq(1-\epsilon)P(A)=\underline{P}(A)$, for all $A\in\mathcal{F}\setminus\mathcal{X}$. In addition, $\check{\Pi}(\mathcal{X})=1>1-\epsilon=\underline{P}(\mathcal{X})$, and so $\check{\Pi}\in\mathcal{M}(\underline{P})$. This shows that $\mathcal{M}(\underline{P}^{\prime})\subseteq\mathcal{M}(\underline{P})$.
For the converse, pick any $\tilde{\Pi}\in\mathcal{M}(\underline{P})$. By definition, we have that $\tilde{\Pi}(A)\geq(1-\epsilon)P(A)=\underline{P}^{\prime}(A)$, for all $A\in\mathcal{F}\setminus\mathcal{X}$. In addition, $\tilde{\Pi}(\mathcal{X})=1\geq 1=\underline{P}^{\prime}(\mathcal{X})$, and so $\tilde{\Pi}\in\mathcal{M}(\underline{P}^{\prime})$. This shows that $\mathcal{M}(\underline{P})\subseteq\mathcal{M}(\underline{P}^{\prime})$.
In turn, we have shown that $\mathcal{M}(\underline{P}^{\prime})=\mathcal{M}(\underline{P})$, and so $\mathcal{M}(\underline{P})$ is a well-defined weak-compact set, and $\underline{P}$ is a well-defined lower probability by Cerreia-Vioglio et al. (2015, Section 2.1.(viii)). ∎
The intuition behind Lemma 6 is that $\underline{P}^{\prime}$ and $\underline{P}$ only disagree (by $\epsilon$ much) on the value to assign to $\mathcal{X}$. But any countably additive probability measure $\Pi$ assigns probability $1$ to the whole state space $\mathcal{X}$. So, to determine whether $\Pi$ belongs to $\mathcal{M}(\underline{P}^{\prime})=\mathcal{M}(\underline{P})$, it is enough to check if $\Pi$ set-wise dominates $\underline{P}^{\prime}$ and $\underline{P}$ on the events in $\mathcal{F}\setminus\mathcal{X}$. An immediate consequence of this argument is that we can write
$\mathcal{M}(\underline{P}^{\prime})=\mathcal{M}(\underline{P})=\{\Pi\in\Delta_{\mathcal{X}}:\Pi(A)\geq\underline{P}(A)=\underline{P}^{\prime}(A)\text{, }\forall A\in\mathcal{F}\setminus\mathcal{X}\}.$
Now, let $\mathcal{Q}_{\epsilon}\subset\Delta_{\mathcal{Y}}$ be a credal set defined similarly to $\mathcal{P}_{\epsilon}$, and consider its “associated” lower probability $\underline{Q}=(1-\epsilon)Q$. We are interested in the Optimal Transport (OT) map between $\mathcal{M}(\underline{P})$ and $\mathcal{M}(\underline{Q})$. Because these sets are completely characterized by $\underline{P}$ and $\underline{Q}$, respectively (as we have seen in Section 2), we focus our attention on such lower probabilities.
@@
Suppose $\mathcal{X}$ and $\mathcal{Y}$ are separable, so that the elements of $\Delta_{\mathcal{X}}$ and $\Delta_{\mathcal{Y}}$ are Radon measures. Let $c:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}_{+}$ be a Borel-measurable (cost) function. Given lower probabilities $\underline{P}$ and $\underline{Q}$ on $\mathcal{X}$ and $\mathcal{Y}$, respectively, we want to find the (measurable) optimal transport map $T:\mathcal{X}\rightarrow\mathcal{Y}$ that attains the infimum
$\inf\left\{{\int_{\mathcal{X}}c(x,T(x))\underline{P}(\text{d}x):T_{\#}\underline{P}=\underline{Q}}\right\}.$ (7)
In Definition 7, we assume that function $x\mapsto c(x,T(x))$ is Choquet integrable with respect to $\underline{P}$ (Definition 3, Proposition 4). Notice that the OT map $T$ need not exist, [^4] As pointed out in Section 1, in the classical (precise) case, an optimal transport map $T$ does not exist when $P$ is a Dirac measure, but $Q$ is not. so we will need to verify its existence in every application of interest. We now show that, for $\epsilon$-contaminated credal sets $\mathcal{P}_{\epsilon}$ and $\mathcal{Q}_{\epsilon}$, LPM is equivalent to the classical Monge’s OT problem of finding the OT map between the contaminated probabilities $P$ and $Q$.
@@
If $\underline{P}$ and $\underline{Q}$ are the lower envelopes of the $\epsilon$-contaminations $\mathcal{P}_{\epsilon}\subseteq\Delta_{\mathcal{X}}$ and $\mathcal{Q}_{\epsilon}\subseteq\Delta_{\mathcal{Y}}$ of $P\in\Delta_{\mathcal{X}}$ and $Q\in\Delta_{\mathcal{Y}}$, respectively, then the LPM of Definition 7 is equivalent to the classical Monge’s OT Problem involving $P$ and $Q$.
@@
By Proposition 4 and (Wasserman and Kadane, 1990, Example 3), we have that
@@
$=\int_{0}^{\infty}\underline{P}\left(\{x\in\mathcal{X}:c(x,T(x))\geq t\}\right)\text{d}t+\int_{-\infty}^{0}\left[\underline{P}\left(\{x\in\mathcal{X}:c(x,T(x))\geq t\}\right)-\underline{P}(\mathcal{X})\right]\text{d}t$
$=\int_{0}^{\infty}(1-\epsilon){P}\left(\{x\in\mathcal{X}:c(x,T(x))\geq t\}\right)\text{d}t$
$+\int_{-\infty}^{0}\left[(1-\epsilon){P}\left(\{x\in\mathcal{X}:c(x,T(x))\geq t\}\right)-(1-\epsilon){P}(\mathcal{X})\right]\text{d}t$
$=(1-\epsilon)\int_{0}^{\infty}{P}\left(\{x\in\mathcal{X}:c(x,T(x))\geq t\}\right)\text{d}t+(1-\epsilon)\int_{-\infty}^{0}\left[{P}\left(\{x\in\mathcal{X}:c(x,T(x))\geq t\}\right)-{P}(\mathcal{X})\right]\text{d}t$
$=(1-\epsilon)\left\{{\int_{0}^{\infty}{P}\left(\{x\in\mathcal{X}:c(x,T(x))\geq t\}\right)\text{d}t+\int_{-\infty}^{0}\left[{P}\left(\{x\in\mathcal{X}:c(x,T(x))\geq t\}\right)-{P}(\mathcal{X})\right]\text{d}t}\right\}$
@@
$T_{\#}\underline{P}(B)=\underline{P}(T^{-1}(B))=(1-\epsilon)P(T^{-1}(B)),\quad\forall B\in\mathcal{G}$ (8)
and
@@
$T_{\#}\underline{P}=\underline{Q}\iff(1-\epsilon)P\circ T^{-1}=(1-\epsilon){Q}\iff P\circ T^{-1}=Q\iff T_{\#}{P}={Q}.$
@@
$\inf\left\{{\int_{\mathcal{X}}c(x,T(x))\underline{P}(\text{d}x):T_{\#}\underline{P}=\underline{Q}}\right\}$
$=\inf\left\{{(1-\epsilon)\int_{\mathcal{X}}c(x,T(x))P(\text{d}x):T_{\#}{P}={Q}}\right\}$
$=\inf\left\{{\int_{\mathcal{X}}c(x,T(x))P(\text{d}x):T_{\#}{P}={Q}}\right\},$ (10)
which is the classical Monge’s OT problem. The last equality holds because $c$ is a positive functional. ∎
To be more precise, Theorem 8 tells us that LPM is equivalent to a slightly more restrictive version of the classical Monge’s OT problem, since we need to require that function $x\mapsto c(x,T(x))$ is Choquet integrable with respect to $\underline{P}$. [^5] If $c$ is bounded, we know that this is the case by Proposition 4. For all practical purposes, though, this does not entail loss of generality, since even for the classical Monge’s OT problem, if we do not assume boundedness of the cost function, we need to check that $\int_{\mathcal{X}}c(x,T(x))P(\text{d}x)$ exists. We now give an example, formulated as a corollary, in which Theorem 8 proves useful.
@@
Let $\mathcal{P}_{\epsilon}$ and $\mathcal{Q}_{\epsilon}$ denote the $\epsilon$-contaminations of countably additive probability measures $P$ and $Q$ on $\mathcal{X}=\mathcal{Y}=\mathbb{R}$. Choose cost function $c$ such that $c(x,y)=h(x-y)$, where $h$ is a strictly convex, positive, Borel-measurable functional. If $P$ and $Q$ have finite $p$-th moment, $p\in[1,\infty)$, and $P$ has no atom, then the unique solution to LPM is $T=F^{-1}_{Q}\circ F_{P},$ where $F_{P}$ and $F_{Q}$ are the cdf’s of $P$ and $Q$, respectively.
@@
3.1 Extension
Adopting the Kantorovich formulation of the OT problem would strengthen our result, since – as we shall see in Corollary 11.1 – a suitable choice of the cost function $c$ will ensure us that the OT map $T$ exists. In addition, since the majority of existing OT results are expressed as a solution of the classical Kantorovich OT problem, we would be able to immediately use them in the context of $\epsilon$-contaminated credal sets.

### v3

@@ 3 Main Results
In this section, we answer the two questions that we put forth in the Introduction. First, we give the general definitions of Monge’s and Kantorovich’s problems for transporting lower probabilities $\underline{P}$ and $\underline{Q}$. Then, we notice how for a special type of cores $\mathcal{M}(\underline{P})$ and $\mathcal{M}(\underline{Q})$, namely those associated with contaminated (countably additive) probabilities $P$ and $Q$, such problems are equivalent to the classical ones, where the transport happens between $P$ and $Q$. This is because $\mathcal{M}(\underline{P})$ and $\mathcal{M}(\underline{Q})$ are completely characterized by the lower probabilities $\underline{P}=(1-\epsilon)P$ and $\underline{Q}=(1-\epsilon)Q$. We show formally how the intuition that the $(1-\epsilon)$ scaling factor does not play a role when solving Monge’s and Kantorovich’s problems is correct. As a result, the lower probabilities and the classical versions of such problems coincide.
We begin by remarking two notational choices and an assumption that we make in the rest of the paper. We will put $\mathcal{M}^{\text{fa}}(\underline{P})\equiv\mathcal{M}(\underline{P})$, and we will call $\Delta^{\text{ca}}_{\mathcal{X}}$ the space of countably additive probabilities on $\mathcal{X}$. We will also assume that the cost function $c$ is (Borel) measurable, so that (i) inner lower probability $\underline{P}^{\star}$ and “classical” lower probability $\underline{P}$ coincide, and (ii) the Choquet integrals that we consider exist.
The special cores that we consider are the so-called $\epsilon$-contaminated credal sets. That is, given a countably additive probability measure $P$ on $\mathcal{X}$, $P\in\Delta^{\text{ca}}_{\mathcal{X}}$, we consider the set
$\begin{split}\mathcal{P}_{\epsilon}=\{\Pi\in\Delta^{\text{fa}}_{\mathcal{X}}:\text{ }&\Pi(A)=(1-\epsilon)P(A)+\epsilon R(A),\\ &\forall R\in\Delta^{\text{fa}}_{\mathcal{X}},\forall A\in\mathcal{F}\},\end{split}$ (3)
where $\epsilon$ is a parameter in $[0,1]$.
@@
Let $\mathcal{P}_{\epsilon}$ be an $\epsilon$-contaminated credal set as in (3). Then, $\underline{P}^{\prime}(A)$ is given by
$\inf_{\Pi\in\mathcal{P}_{\epsilon}}\Pi(A)=\begin{cases}(1-\epsilon)P(A),&\text{for all }A\in\mathcal{F}\setminus\{\mathcal{X}\}\\ 1,&\text{for }A=\mathcal{X}\end{cases}$ (4)
@@
$\mathcal{P}_{\epsilon}=\mathcal{M}(\underline{P}^{\prime})=\{\Pi\in\Delta^{\text{{\em fa}}}_{\mathcal{X}}:\Pi(A)\geq\underline{P}^{\prime}(A)\text{, }\forall A\in\mathcal{F}\}.$ (5)
@@
Both these properties were proven in Wasserman and Kadane (1990, Example 3) and in Walley (1991, Section 2.9.2). ∎
A remark is in order. The elements of $\mathcal{P}_{\epsilon}$ must be finitely additive probabilities, and not merely countably additive, because if that were not the case, then $\mathcal{P}_{\epsilon}$ would not be weak-compact, and so it would not be a well-defined core. Let us give an example, borrowed from Walley (1991). Consider an $\epsilon$-contamination model on the Naturals $\mathbb{N}$, with $\epsilon=1$ (this is the vacuous lower probability that assigns lower probability 00 to every natural number). The sequence $(\delta_{n})_{n\in\mathbb{N}}$ of Dirac measures $\delta_{n}$ assigning mass $1$ to ${n}\in\mathbb{N}$ is a sequence of countably additive probability measures that belongs to $\mathcal{P}_{\epsilon}$. But this sequence has no weak converging subsequence to a countably additive probability measure. If it did, it would have to assign probability 00 to all of the Naturals $n\in\mathbb{N}$. Hence, $\mathcal{P}_{\epsilon}$ cannot be weak-compact in this case.
This is a technicality which does not affect the interpretation of our results, for two main reasons. First, all countably additive probabilities are also finitely additive, that is, $\Delta^{\text{ca}}_{\mathcal{X}}\subset\Delta^{\text{fa}}_{\mathcal{X}}$. Second, we consider contaminations of a countably additive probability $P\in\Delta^{\text{ca}}_{\mathcal{X}}$. [^7] Of course, in general we may have $P\in\Delta^{\text{fa}}_{\mathcal{X}}$. This is because we want to relate the lower probability versions of Monge’s and Kantorovich’s OT problems (that we introduce later in this Section) with the classical ones, that are formulated for countably additive probabilities.
In the remainder of the paper, we will work with the (incoherent, according to Walley (1991, Section 2.5)) lower probability $\underline{P}$ such that
@@
in place of $\underline{P}^{\prime}$. This is an unnormalized countably additive measure, which is Radon measure if the underlying space is separable. The reason we work with $\underline{P}$ in (6) is twofold: calculations are easier to carry out, and also the following lemma holds. The interested reader can find a further discussion on this choice in Appendix A.
@@
Pick any countably additive probability measure $P$ on $\mathcal{X}$, any $\epsilon\in[0,1]$, and consider the two lower probabilities $\underline{P}^{\prime}$ and $\underline{P}$ in (4) and (6), respectively. Let $\mathcal{M}(\underline{P})=\{\Pi\in\Delta^{\text{\em fa}}_{\mathcal{X}}:\Pi(A)\geq\underline{P}(A)\text{, }\forall A\in\mathcal{F}\}$. Then, $\mathcal{M}(\underline{P}^{\prime})=\mathcal{M}(\underline{P})$.
@@
We begin by noting that $\underline{P}^{\prime}$ is a well-defined lower probability by Cerreia-Vioglio et al. (2015, Section 2.1.(viii)). Now, pick any $\Pi\in\Delta^{\text{fa}}_{\mathcal{X}}$. We have that $\Pi(A)\geq\underline{P}(A)$ if and only if $\Pi(A)\geq\underline{P}^{\prime}(A)$, for all $A\in\mathcal{F}\setminus\{\mathcal{X}\}$. In addition, $\Pi(\mathcal{X})=\underline{P}^{\prime}(\mathcal{X})=1>1-\epsilon=\underline{P}(\mathcal{X})$. In turn, this shows that $\mathcal{M}(\underline{P}^{\prime})=\mathcal{M}(\underline{P})$, concluding the proof. ∎
The intuition behind Lemma 6 is that $\underline{P}^{\prime}$ and $\underline{P}$ only disagree (by $\epsilon$ much) on the value to assign to $\mathcal{X}$. But any finitely additive probability measure $\Pi$ assigns probability $1$ to the whole state space $\mathcal{X}$. So, to determine whether $\Pi$ belongs to $\mathcal{M}(\underline{P}^{\prime})=\mathcal{M}(\underline{P})$, it is enough to check if $\Pi$ set-wise dominates $\underline{P}^{\prime}$ and $\underline{P}$ on the events in $\mathcal{F}\setminus\{\mathcal{X}\}$. An immediate consequence of this argument is that we can write $\mathcal{M}(\underline{P}^{\prime})=\mathcal{M}(\underline{P})=\{\Pi\in\Delta^{\text{fa}}_{\mathcal{X}}:\Pi(A)\geq\underline{P}(A)=\underline{P}^{\prime}(A)\text{, }\forall A\in\mathcal{F}\setminus\{\mathcal{X}\}\}.$
Now, let $\mathcal{Q}_{\epsilon}\subset\Delta^{\text{fa}}_{\mathcal{Y}}$ be a credal set defined similarly to $\mathcal{P}_{\epsilon}$, and consider its “associated” lower probability $\underline{Q}=(1-\epsilon)Q$, $Q\in\Delta^{\text{ca}}_{\mathcal{Y}}$. We are interested in the Optimal Transport (OT) map between $\mathcal{M}(\underline{P})$ and $\mathcal{M}(\underline{Q})$. Because these sets are completely characterized by $\underline{P}$ and $\underline{Q}$, respectively (as we have seen in Section 2), we focus our attention on such lower probabilities.
3.1 Lower Probability Monge’s (LPM) Problem
@@
Let $c:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}_{+}$ be a Borel measurable (cost) function. Given lower probabilities $\underline{P}$ and $\underline{Q}$ on $\mathcal{X}$ and $\mathcal{Y}$, respectively, we want to find the (measurable) optimal transport map $T:\mathcal{X}\rightarrow\mathcal{Y}$ that solves the following optimization problem
$\operatorname*{arg\,inf}\left\{{\int_{\mathcal{X}}c(x,T(x))\underline{P}(\text{d}x):T_{\#}\underline{P}=\underline{Q}}\right\}.$ (7)
We assume $c$ to be Borel measurable and $T$ to be measurable to ensure that $c(x,T(x))$ is Choquet integrable. Let us also notice that we work with the Choquet integral $\int_{\mathcal{X}}c(x,T(x))\underline{P}(\text{d}x)$ because, in the case of $\epsilon$-contaminations, it corresponds to the lower expectation $\inf_{P\in\mathcal{M}(\underline{P})}\int_{\mathcal{X}}c(x,T(x))P(\text{d}x)$; this is because $\underline{P}$ is $2$-monotone. In general, though, we have that $\inf_{P\in\mathcal{M}(\underline{P})}\int_{\mathcal{X}}c(x,T(x))P(\text{d}x)\leq\int_{\mathcal{X}}c(x,T(x))\underline{P}(\text{d}x)$; the Choquet integral (with respect to a lower probability) is an upper bound for the worst-case (i.e. lower) expectation. In the future, we will study how the solution to LPM changes if we consider more general cores (i.e. not necessarily $\epsilon$-contaminations), and if we work with lower expectations in place of Choquet integrals.
Notice that the OT map $T$ need not exist, [^8] In the classical (precise) case, an optimal transport map $T$ does not exist when $P$ is a Dirac measure, but $Q$ is not. so we will need to verify its existence in every application of interest. We now show that, for $\epsilon$-contaminated credal sets $\mathcal{P}_{\epsilon}$ and $\mathcal{Q}_{\epsilon}$, LPM is equivalent to the classical Monge’s problem of finding the OT map between the contaminated probabilities $P$ and $Q$. Throughout the rest of the paper, we perpetrate an abuse of terminology and refer to $\underline{P}$ and $\underline{Q}$ as “lower envelopes”, even though they are not normalized at $1$.
@@
Suppose $\mathcal{X}$ and $\mathcal{Y}$ are separable, so that the elements of $\Delta^{\text{\em ca}}_{\mathcal{X}}$ and $\Delta^{\text{\em ca}}_{\mathcal{Y}}$ are Radon measures. If $\underline{P}$ and $\underline{Q}$ are the lower envelopes of the $\epsilon$-contaminations $\mathcal{P}_{\epsilon}\subseteq\Delta^{\text{\em fa}}_{\mathcal{X}}$ and $\mathcal{Q}_{\epsilon}\subseteq\Delta^{\text{\em fa}}_{\mathcal{Y}}$ of $P\in\Delta^{\text{\em ca}}_{\mathcal{X}}$ and $Q\in\Delta^{\text{\em ca}}_{\mathcal{Y}}$, respectively, then the LPM of Definition 7 is equivalent to the classical Monge’s OT Problem involving $P$ and $Q$.
@@
By Proposition 4, Corollary 4.1, and Lemma 6, we have that
@@
$=\int_{0}^{\infty}\underline{P}\left(\{x\in\mathcal{X}:c(x,T(x))\geq t\}\right)\text{d}t$
$=\int_{0}^{\infty}(1-\epsilon){P}\left(\{x\in\mathcal{X}:c(x,T(x))\geq t\}\right)\text{d}t$ (8)
$=(1-\epsilon)\int_{0}^{\infty}{P}\left(\{x\in\mathcal{X}:c(x,T(x))\geq t\}\right)\text{d}t$
@@
$T_{\#}\underline{P}(B)=\underline{P}(T^{-1}(B))=(1-\epsilon)P(T^{-1}(B)),$ (9)
for all $B\in\mathcal{G}$, and
@@
$T_{\#}\underline{P}=\underline{Q}$ $\iff(1-\epsilon)P\circ T^{-1}=(1-\epsilon){Q}$
$\iff P\circ T^{-1}=Q\iff T_{\#}{P}={Q}.$
@@
$\operatorname*{arg\,inf}\left\{{\int_{\mathcal{X}}c(x,T(x))\underline{P}(\text{d}x):T_{\#}\underline{P}=\underline{Q}}\right\}$
$=\operatorname*{arg\,inf}\left\{{(1-\epsilon)\int_{\mathcal{X}}c(x,T(x))P(\text{d}x):T_{\#}{P}={Q}}\right\}$
$=\operatorname*{arg\,inf}\left\{{\int_{\mathcal{X}}c(x,T(x))P(\text{d}x):T_{\#}{P}={Q}}\right\},$ (11)
which is the classical Monge’s OT problem. The last equality holds because $c$ is a positive functional. Notice also that since $\underline{P}$ and $\underline{Q}$ are both countably additive measures normed to $(1-\epsilon)$, the Choquet integrals in this proof coincide with classical Lebesgue-Stieltjes integrals. ∎
In Theorem 8 we assume separability of $\mathcal{X}$ and $\mathcal{Y}$ to further relate the LPM and the classical Monge’s OT problems. Indeed, in the latter, separability ensures the existence of Borel measurable selections, which is crucial for defining transport maps. Notice also that for Theorem 8 to hold we do not need to implicitly assume that the contaminating parameter $\epsilon$ is the same for both $\mathcal{P}_{\epsilon}$ and $\mathcal{Q}_{\epsilon}$. That is, we could consider $\mathcal{P}_{\epsilon}$ and $\mathcal{Q}_{\epsilon^{\prime}}$, $\epsilon^{\prime}\neq\epsilon$. This because, for the equivalences below (10) to work, it must be that $(1-\epsilon)/(1-\epsilon^{\prime})=1$, and so $\epsilon^{\prime}=\epsilon$ must hold.
@@
Let $\mathcal{P}_{\epsilon}$ and $\mathcal{Q}_{\epsilon}$ denote the $\epsilon$-contaminations of countably additive probability measures $P$ and $Q$ on $\mathcal{X}=\mathcal{Y}=\mathbb{R}$. Choose cost function $c$ such that $c(x,y)=h(x-y)$, where $h$ is a strictly convex, positive, Borel measurable functional. If $P$ and $Q$ have finite $p$-th moment, $p\in[1,\infty)$, and $P$ has no atom, then the unique solution to LPM is $T=F^{-1}_{Q}\circ F_{P},$ where $F_{P}$ and $F_{Q}$ are the cdf’s of $P$ and $Q$, respectively.
@@
3.2 Lower Probability Kantorovich’s (LPK) Problem
Adopting the Kantorovich formulation of the OT problem would strengthen our result, since – as we shall see in Corollary 13.1 – a suitable choice of the cost function $c$ will ensure us that the OT map $T$ exists. In addition, since most existing OT results are expressed as a solution to the classical Kantorovich OT problem, we would be able to immediately use them in the context of $\epsilon$-contaminated credal sets.